Help me please
2x+5y=3
-x+3y=-7

Answer by checkley71(8403)

-X+3Y=-7 OR -X=-3Y-7 OR X=3Y+7 NOW REPLACE X IN THE OTHERE EQUATION BY (3Y+7) & SOLVE FOR Y
2(3Y+7)+5Y=3
6Y+14+5Y=3
11Y=3-14
11Y=-11
Y=-11/11
Y=-1 ANSWER.
-X+3*-1=-7
-X-3=-7
-X=-7+3
-X=-4
X=4 ANSWER.